A WET OPEN UMBRELLA IS HELD VERTICAL AND IT IS WHIRLED ABOUT THE HANDLE AT A UNIFORM RATE OF 21 REV. IN 44s. IF THE RIM OF THE UMBRELLA IS A CIRCLE OF 1m IN DIAMETER AND THE HEIGHT OF THE RIM ABOVE THE FLOOR IS 4.9m. THE LOCUS OF THE DROP IS A CIRCLE OF RADIUS :
1. (2.5)^1/2 m
2. 1 m
3. 3 m
4. 1.5 m
1
Expert's answer
2014-03-31T12:38:45-0400
Find the time of drop falling:
g* (t^2) /2 = 4.9
t = sqrt(2 *4.9 /g)
Find the normal accelration:
w = 21 * 2 * pi / 44
an = R * w^2 = 9
Find the circle radius:
R = 1 + an * (t^2) /2 = 1 + an * 4.9 /g = 5.5 (m)